OpenStudy (anonymous):

8x^2-2x-15=0 How do I solve this using quadratic formula ?

OpenStudy (anonymous):

the quadratic formula is +-b squareroot of b^2-4ac all over 2a. so what you do is you use 8=a, -2=b and -15=c. plug in.

OpenStudy (anonymous):

use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=8,b=-2,c=-15\]

OpenStudy (saifoo.khan):

HELPERS!

OpenStudy (anonymous):

nusrat fateh saifoo khan is on over drive

OpenStudy (anonymous):

Aw! I wanted to use my saved copy of the quadratic.....

OpenStudy (anonymous):

- They say its 2 fractions ? Im confused .

OpenStudy (saifoo.khan):

@satellite haha, from where did u got that name? lol

OpenStudy (anonymous):

that's the plus/minus bit in the middle...

OpenStudy (anonymous):

See what I mean? U will end up with 2 answers...

OpenStudy (anonymous):

when you have the +-, one answer will use the positive and one will use the negative.

OpenStudy (anonymous):

by some miracle this actually factors sa \[(2x-3)(4x+5)=0\] so you know \[b^2-4ac\] will be a perfect square and your answers will be \[x=\frac{3}{2},x=-\frac{5}{4}\]

OpenStudy (anonymous):

write it with equation forum

OpenStudy (anonymous):

(x+3)/(x+2) = 6/5 do cross multiplication 6(x+2) + 5(x+3) 6x + 12 = 5x + 15 transpose 5x to left hand side and 12 to right hand side 6x-5x = 15-12 x = 3